OpenStudy (anonymous):

I have these two points, (7,3) and (8,5). I solved on Mathway and got y = 2x + 11. Can someone explain how I can find that?

OpenStudy (anonymous):

@thomaster

OpenStudy (anonymous):

@Hero

hero (hero):

Explain what you did to solve it on mathway.

OpenStudy (anonymous):

I gave the engine the two points, and pressed "Find the equation using Point-Slope"

OpenStudy (anonymous):

I've been stuck on this problem for a while and resulted to this

OpenStudy (anonymous):

I solved my first problem using point - slope, and it worked fine, I'm not sure why this one is giving me trouble.

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ &({\color{red}{ 7}}\quad ,&{\color{blue}{ 3}})\quad &({\color{red}{ 8}}\quad ,&{\color{blue}{ 5}}) \end{array} \\\quad \\ slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (anonymous):

I know, I have the slope already

OpenStudy (anonymous):

The slope is 2

OpenStudy (jdoe0001):

ok... so plug it in and solve for "y" then :)

OpenStudy (anonymous):

That's what I'm having trouble with

OpenStudy (anonymous):

I haven't been getting y = 2x + 11

OpenStudy (anonymous):

Which is why I've come to you guys for help

OpenStudy (jdoe0001):

that is correct \(\bf y-{\color{blue}{ 3}}={\color{green}{ 2}}(x-{\color{red}{ 7}})\implies y-3=2x-14\implies y=2x-14+3\)

OpenStudy (anonymous):

Wouldn't you get y = 2x - 17?

hero (hero):

@jdoe0001 he solved it on mathway but couldn't figure out how to solve it on his own.

OpenStudy (jdoe0001):

well... "he" didn't "solve" it in mathway mathway simply gave the answer heehe

OpenStudy (anonymous):

I do know how to solve it, but I just can't understand why I'm getting it wrong

OpenStudy (anonymous):

I use mathway sometimes to check my answer when I'm uncertain of my own

OpenStudy (jdoe0001):

well..... look at => \(\bf y-{\color{blue}{ 3}}={\color{green}{ 2}}(x-{\color{red}{ 7}})\implies y-3=2x-14\implies y=2x-14+3\) anything you find wrong there?

OpenStudy (jdoe0001):

superhelp101 that sounds fair :)

OpenStudy (anonymous):

:>

OpenStudy (anonymous):

But no, I don't see anything wrong

OpenStudy (anonymous):

Oh, wait a second

OpenStudy (anonymous):

Nevermind

OpenStudy (anonymous):

lol

OpenStudy (jdoe0001):

I think you got -17 because you didn't add \(\bf +3\) you subtracted \(\bf -3\) to both sides and to eliminate it from the left side, you have to add it

OpenStudy (anonymous):

No, I added it to both sides

OpenStudy (anonymous):

Let me show you how I've been solving it

OpenStudy (jdoe0001):

k

OpenStudy (anonymous):

y - 3 = 2(x - 7) y + 3= 2(x - 7 + 3) y = 2x - 10

OpenStudy (anonymous):

Dangit, forgot to do the multiplication

OpenStudy (anonymous):

y = 2x - 17***

OpenStudy (jdoe0001):

?

OpenStudy (anonymous):

The second one is my answer

OpenStudy (jdoe0001):

y - 3 = 2(x - 7) y + 3= 2(x - 7 + 3) <------ \(\Large ?\)

OpenStudy (jdoe0001):

yes.. but how did you get it?

OpenStudy (anonymous):

I showed up above, but here it is again. y - 3 = 2(x - 7) y + 3= 2(x - 7 + 3) y = 2x - 10

OpenStudy (radar):

You see I don't get it either.